OpenStudy (anonymous):

Help Please!(: Consider the function f(x) = 6sqrt {x}+7 on the interval [ 1 , 10 ]. Find the average or mean slope of the function on this interval.

OpenStudy (kainui):

Well if you want to find the average, what do you do normally? Don't even think calculus yet.

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

Nope, the average is like when you have 3 people and each person has their own height, like 5 feet tall, 6 feet tall, and 4 feet tall. So you add up all the heights and divide by the total number, So what's the average height of the people?

OpenStudy (anonymous):

5 feet @Kainui

OpenStudy (kainui):

Alright, cool. So now you need to find the formula for slope of that function. Then this is what you're taking the average of. You need to add up all the individual amounts of slope over that function from the boundary points (1 to 10) by integration and divide by the total interval which is (10-1). It's basically the same as the average.

OpenStudy (anonymous):

I don't know how to go about starting this problem.. add of the individual amounts of the slope over the function? how do i do that? @kainui

OpenStudy (dumbcow):

the integral definition of avg slope is: \[\frac{1}{b-a}\int\limits_{a}^{b}f'(x) dx\] using fund thm of calculus, this reduces to \[\frac{f(b)-f(a)}{b-a}\] which is the same as simple slope formula for the 2 endpoints of interval

OpenStudy (dumbcow):

so by drawing a line from the endpoints, the slope of that line is no different than if you averaged all the slopes at every point on the curve

OpenStudy (anonymous):

So how do I go about doing this problem. I understand i need to find the slope. but how?

OpenStudy (dumbcow):

from algebra 1 you learned slope is change in y over change in x

OpenStudy (kainui):

That's what the derivative is all about. Remember y=mx+b, where m is the slope? That's rise over run, which is (y2-y1)/(x2-x1) which is essentially what the definition of a derivative is, but it's [f(x+h)-f(x)]/[(x+h)-(x)] as h approaches 0, so it's instantaneous.

OpenStudy (anonymous):

yes but there's only one point. (1,10)

OpenStudy (dumbcow):

no thats the interval...x varies from 1 to 10 plug those into function to get y values

OpenStudy (dumbcow):

wow you prob didn't get anything i wrote huh

OpenStudy (anonymous):

yeahh... I'm so confused int his class... But i'm trying. so 13 and 25.97? haha

OpenStudy (dumbcow):

yes..i would leave it in exact form though

OpenStudy (kainui):

Well pdd what you have to understand is that an integral is the reverse operation of a derivative, so you are kind of not calculating anything much since you're basically doing something, undoing it, and then plugging in those values... lol

OpenStudy (anonymous):

so... 6sqrt{1}7 and 6sqrt{10}+7 ?

OpenStudy (dumbcow):

no need for sqrt1 ... 13 and 6sqrt10 +7 now plug those into slope formula and simplify

OpenStudy (dumbcow):

\[\rightarrow \frac{(6\sqrt{10}+7) -13}{10-1} = \frac{6\sqrt{10}-6}{9} = \frac{2(\sqrt{10}-1)}{3}\]

OpenStudy (anonymous):

I was just about to type that out.... but for {1,10}

OpenStudy (anonymous):

By the Mean Value Theorem, we know there exists a c in the open interval ( 1 , 10 ) such that f'( c) is equal to this mean slope. For this problem, there is only one c that works. how do I find it?

OpenStudy (anonymous):

i find the derivative of the function first right?

OpenStudy (dumbcow):

right set f'(x) equal to avg slope found above ... solve for "x" and that is your c value

OpenStudy (anonymous):

so f'(x)=3/sqrt(x) then x becomes c. so 3/sqrt(c)=2(sqrt(10)-1)? and c=0.230886157?

OpenStudy (dumbcow):

derivative is right .... the "c" value seems way too small wait it has to be bigger than 1